The length of a rectangle is 5inlonger than its width. If the perimeter of the rectangle is 46 in, find its area.





Answer by macston(5194)

.
W=width; L=length=W+5in; P=perimeter=46 inches; A=area
.
P=2(L+W) Substitute for L.
46in=2(W+5in+W) Divide each side by 2.
23in=2W+5in Subtract 5in from each side.
18in=2W Divide each side by 2.
9in=W The width is 9 inches.
L=W+5in=9in+5in=14in The length is 14 inches.
.
A=LW
A=(14in)(9in)
A=126 square inches
ANSWER: The area is 126 square inches.